OpenStudy (anonymous):

The sum of two numbers is 51. Their difference is 13. What is the smaller of the two numbers? A. 18 B. 19 C. 20 D. 21 The sum of two numbers is 51. Their difference is 13. What is the smaller of the two numbers? A. 18 B. 19 C. 20 D. 21 @Mathematics

OpenStudy (anonymous):

x+y=51 x-y=13 2x=64 x=32 y=51-32 = 19 so its B

OpenStudy (anonymous):

X=3, Y=5